### v1

@@ Abstract
Multi-task dense scene understanding, which learns a model for multiple dense prediction tasks, has a wide range of application scenarios. Modeling long-range dependency and enhancing cross-task interactions are crucial to multi-task dense prediction. In this paper, we propose MTMamba, a novel Mamba-based architecture for multi-task scene understanding. It contains two types of core blocks: self-task Mamba (STM) block and cross-task Mamba (CTM) block. STM handles long-range dependency by leveraging Mamba, while CTM explicitly models task interactions to facilitate information exchange across tasks. Experiments on NYUDv2 and PASCAL-Context datasets demonstrate the superior performance of MTMamba over Transformer-based and CNN-based methods. Notably, on the PASCAL-Context dataset, MTMamba achieves improvements of +2.08, +5.01, and +4.90 over the previous best method in the tasks of semantic segmentation, human parsing, and object boundary detection, respectively. The code is available at https://github.com/EnVision-Research/MTMamba.
@@ 1 Introduction
Multi-task dense scene understanding is an essential problem in computer vision [36] and has a variety of practical applications, such as autonomous driving [20, 23], healthcare [19], and robotics [48]. It aims to train a model for simultaneously handling multiple dense prediction tasks, such as semantic segmentation, monocular depth estimation, surface normal estimation, and object boundary detection.
@@ 2.1 Multi-Task Learning
Multi-task learning (MTL) is a learning paradigm that aims to simultaneously learn multiple tasks in a single model [50]. Recent MTL research mainly focuses on multi-objective optimization [25, 24, 45, 44, 34, 47, 26] and network architecture design [46, 43, 42, 37]. In multi-task dense scene understanding, most existing works focus on designing architecture [36], especially designing specific modules in the decoder to achieve better cross-task interaction. For example, based on CNN, Xu et al. [42] introduce PAD-Net, incorporating an effective multi-modal distillation module to promote information fusion between different tasks in the decoder. MTI-Net [37] is a complex multi-scale and multi-task CNN architecture with an information distillation across various feature scales. As the convolution operation mainly captures local features [2], recent approaches [46, 43] utilize the attention mechanism to grasp global context and develop Transformer-based decoders for multi-task scene understanding. For instance, Ye & Xu [46] introduce InvPT, a Transformer-based multi-task architecture, employing an effective UP-Transformer block for multi-task feature interaction at different feature scales. MQTransformer [43] designs a cross-task query attention module to enable effective task association and information exchange in the decoder.
@@ 2.2 State Space Models
Many different structures have been proposed recently to improve the expressivity and efficiency of SSMs. Gu et al. [14] propose structured state space models (S4) to improve computational efficiency, where the state matrix is a sum of low-rank and normal matrices. Many follow-up works attempt to enhance the effectiveness of S4. For example, Fu et al. [11] design a new SSM layer H3 to fill the performance gap between SSMs and Transformers in language modeling. Mehta et al. [32] introduce a gated state space layer using gated units for improving expressivity.
Recently, Gu & Dao [13] further propose Mamba with the core operation S6, an input-dependent selection mechanism of S4, which achieves linear scaling in sequence length and demonstrates superior performance over Transformers on various benchmarks. Mamba has been successfully applied in image classification [27, 53], image segmentation [41], and graph prediction [38]. Different from them, which use Mamba in the single-task setting, we consider a more challenging multi-task setting and propose novel self-task and cross-task Mamba modules to capture intra-task and inter-task dependence.
@@ 3.4 Mamba-based Decoder
Mamba-based Feature Extractor (MFE). We introduce a Mamba-based feature extractor to learn the representation of 2D images. It is a critical module in the proposed Mamba-based decoder. As shown in Figure 2(a), motivated by [13], MFE consists of a linear layer used to expand the feature dimension by a controllable expansion factor $\alpha$, a convolution layer with an activation function for extracting local features, an SS2D operation for modeling long-range dependency, and a layer normalization to normalize the learned features. More formally, given the input feature ${\bf z}$, the output $\bar{{\bf z}}$ of MFE is calculated as
@@
$\tilde{{\bf z}}$ $=\texttt{MFE}({\bf z}_{\text{LN}}),\quad{\bf g}=\sigma(\texttt{Linear}({\bf z}_{\text{LN}})),$ (15)
@@ 4.1 Experimental Setups
Following [46, 43], experiments are conducted on two widely used benchmark datasets with multi-task labels: NYUDv2 [35] and PASCAL-Context [6]. The NYUDv2 dataset comprises a variety of indoor scenes, containing 795 and 654 RGB images for training and testing, respectively. It consists of four tasks: $40$-class semantic segmentation (Semseg), monocular depth estimation (Depth), surface normal estimation (Normal), and object boundary detection (Boundary). The PASCAL-Context dataset, derived from the PASCAL dataset [9], includes both indoor and outdoor scenes and provides pixel-wise labels for tasks like semantic segmentation, human parsing (Parsing), and object boundary detection, with additional labels for surface normal estimation and saliency detection tasks generated by [31]. It contains 4,998 training images and 5,105 testing images.
@@
Table 1: Comparison with state-of-the-art methods on the NYUDv2 dataset. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance. The best and second best results are highlighted in bold and underline, respectively.
@@
Evaluation Metrics.
Table 2: Comparison with state-of-the-art methods on the PASCAL-Context dataset. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance. The best and second best results are highlighted in bold and underline, respectively.
@@ 4.2 Comparison with State-of-the-art Methods
Table 1 shows the comparison on the NYUDv2 dataset. As can be seen, the proposed MTMamba method shows superior performance in all four tasks. For example, the performance of the semantic segmentation task has notably improved from the Transformer-based methods (i.e., InvPT and MQTransformer), increasing by +2.26 and +0.98, respectively, which demonstrates the effectiveness of MTMamba. A qualitative comparison with InvPT is shown in Figure 5, which shows that MTMamba generates more accurate predictions.
Table 2 provides the comparison on the PASCAL-Context dataset. These results show the clear superiority of the proposed MTMamba method. Especially, MTMamba significantly improves the previous best by +2.08, +5.01, and +4.90 in semantic segmentation, human parsing, and object boundary detection tasks, respectively, showing the effectiveness of MTMamba again. Also, a qualitative comparison with InvPT is shown in Figure 4, showing that MTMmaba provides more precise predictions and details.
4.3 Model Analysis
Effectiveness of STM and CTM blocks.
The proposed MTMamba contains two types of core blocks: STM and CTM blocks. We perform an experiment on the NYUDv2 dataset to study the effectiveness of each block with a Swin-Large Transformer encoder. The results are shown in Table 3. “Swin only” (resp. “STM only”) denotes each task only uses two specific Swin Transformer (resp. STM) blocks in each decoder stage. “Single-task” is the single-task counterpart of “Swin only”, which indicates each task has its task-specific model. “STM+CTM” is the default method of MTMamba, i.e., adding a shared CTM block in each decoder stage compared with “STM only”.
Table 3: Effectiveness of the proposed STM and CTM blocks on the NYUDv2 dataset. Swin-Large encoder is used in this experiment. “Swin only” (or “STM only”) denotes each task only uses two specific Swin Transformer (or STM) blocks in each decoder stage. “Single-task” is the single-task counterparts of “Swin only”. “STM+CTM” is the default method of MTMamba. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance.
Method Semseg Depth Normal Boundary $\Delta_{m}$[%] #Param FLOPs
@@ 4.3 Model Analysis
Single-task 54.32 0.5166 19.21 77.30 +0.00 888.77 1074.79
Swin only 53.72 0.5239 19.97 76.50 -1.87 303.18 466.35
STM only 54.66 0.4984 18.81 78.20 +1.84 276.48 435.47
STM+CTM 55.82 0.5066 18.63 78.70 +2.38 307.99 540.81
According to Table 3, “STM only” largely outperforms “Swin only”, demonstrating that the STM block is more effective than the Swin Transformer block. Furthermore, “STM only” has fewer parameters and FLOPs than “Swin only”, which shows the efficiency of the STM block. Compared with “STM only”, “STM+CTM” performs better, confirming the CTM block is beneficial. Moreover, the default configuration (i.e., “STM+CTM”) significantly outperforms “Single-task” on all tasks, demonstrating the effectiveness of MTMamba.
Table 4: Effectiveness of MFE module in MTMamba on the NYUDv2 dataset. Swin-Large encoder is used in this experiment. “W-MSA” is the window-based multi-head self-attention module in Swin Transformer. “MFE” denotes all MFE modules in both STM and CTM blocks.
@@
Table 5: Effectiveness of linear gate in MTMamba on the NYUDv2 dataset. Swin-Large encoder is used in this experiment. “W-MSA” is the window-based multi-head self-attention module in Swin Transformer. “Linear” denotes all linear gates in both STM and CTM blocks.
@@
As shown in Figure 2, the MFE module is SSM-based and is the core of both STM and CTM blocks. We conduct an experiment by replacing all MFE modules in MTMamba with the attention module on the NYUDv2 dataset. As shown in Table 4, MFE is more effective and efficient than attention.
@@
The core of the CTM block is the cross-task interaction, i.e., Equation (23), where we fuse task-specific representation $\tilde{{\bf z}}^{t}$ and shared representation $\tilde{{\bf z}}^{\text{sh}}$ via a task-specific gate ${\bf g}^{t}$. In this experiment, we study its effectiveness by comparing it with the cases of ${\bf g}^{t}=0$ and ${\bf g}^{t}=1$. The experiments are conducted with a Swin-Large Transformer encoder on the NYUDv2 dataset. The results are shown in Table 6. As can be seen, using a specific $\tilde{{\bf z}}^{t}$ (i.e., the case of ${\bf g}^{t}=0$) or shared $\tilde{{\bf z}}^{\text{sh}}$ (i.e., the case of ${\bf g}^{t}=1$) degrades the performance, demonstrating that the adaptive fusion is better.
@@ 4.4 Qualitative Evaluations
Figure 3 shows the comparison of the final decoder feature between the proposed MTMamba and Transformer-based method InvPT [46] in the semantic segmentation task. As can be seen, our method highly activates the regions with contextual and semantic information, which means it captures more discriminative features, resulting in better segmentation performance.
@@
We conduct qualitative studies by comparing the output predictions from our proposed MTMamba against the state-of-the-art Transformer-based method, InvPT [46]. Figures 5 and 4 show the qualitative results on the NYUDv2 and PASCAL-Context datasets, respectively. As can be seen, our method has better visual results than InvPT in all tasks. For example, as highlighted with yellow circles in Figure 5, MTMamba generates more accurate results with better alignments for the semantic segmentation task and clearer object boundaries for the object boundary detection task. Figure 4 demonstrates that MTMamba produces better predictions with more accurate details (like the fingers as highlighted) for both semantic segmentation and human parsing tasks and more distinct boundaries for the object boundary detection task. Hence, both qualitative study (Figures 5 and 4) and quantitative study (Tables 1 and 2) show the superior performance of the proposed MTMamba method.
@@ 5 Conclusion
In this paper, we propose MTMamba, a novel multi-task architecture with a Mamba-based decoder for multi-task dense scene understanding. With two novel blocks (STM and CTM blocks), MTMamba can effectively model long-range dependency and achieve cross-task interaction. Experiments on two benchmark datasets demonstrate that the proposed MTMamba achieves better performance than previous CNN-based and Transformer-based methods.

### v2

@@ Abstract
Multi-task dense scene understanding, which learns a model for multiple dense prediction tasks, has a wide range of application scenarios. Modeling long-range dependency and enhancing cross-task interactions are crucial to multi-task dense prediction. In this paper, we propose MTMamba, a novel Mamba-based architecture for multi-task scene understanding. It contains two types of core blocks: self-task Mamba (STM) block and cross-task Mamba (CTM) block. STM handles long-range dependency by leveraging Mamba, while CTM explicitly models task interactions to facilitate information exchange across tasks. Experiments on NYUDv2 and PASCAL-Context datasets demonstrate the superior performance of MTMamba over Transformer-based and CNN-based methods. Notably, on the PASCAL-Context dataset, MTMamba achieves improvements of +2.08, +5.01, and +4.90 over the previous best methods in the tasks of semantic segmentation, human parsing, and object boundary detection, respectively. The code is available at https://github.com/EnVision-Research/MTMamba.
@@ 1 Introduction
Multi-task dense scene understanding is an essential problem in computer vision [36] and has a variety of practical applications, such as autonomous driving [20, 23], healthcare [19], and robotics [49]. It aims to train a model for simultaneously handling multiple dense prediction tasks, such as semantic segmentation, monocular depth estimation, and surface normal estimation.
@@ 2.1 Multi-Task Learning
Multi-task learning (MTL) is a learning paradigm that aims to simultaneously learn multiple tasks in a single model [51]. Recent MTL research mainly focuses on multi-objective optimization [25, 24, 45, 44, 34, 48, 26, 46] and network architecture design [47, 43, 42, 37]. In multi-task dense scene understanding, most existing works focus on designing architecture [36], especially designing specific modules in the decoder to achieve better cross-task interaction. For example, based on CNN, Xu et al. [42] introduce PAD-Net, incorporating an effective multi-modal distillation module to promote information fusion between different tasks in the decoder. MTI-Net [37] is a complex multi-scale and multi-task CNN architecture with an information distillation across various feature scales. As the convolution operation mainly captures local features [2], recent approaches [47, 43] utilize the attention mechanism to grasp global context and develop Transformer-based decoders for multi-task scene understanding. For instance, Ye & Xu [47] introduce InvPT, a Transformer-based multi-task architecture, employing an effective UP-Transformer block for multi-task feature interaction at different feature scales. MQTransformer [43] designs a cross-task query attention module to enable effective task association and information exchange in the decoder.
@@ 2.2 State Space Models
Many different structures have been proposed recently to improve the expressivity and efficiency of SSMs. Gu et al. [14] propose structured state space models (S4) to improve computational efficiency, where the state matrix is a sum of low-rank and normal matrices. Many follow-up works attempt to enhance the effectiveness of S4. For example, Fu et al. [11] design a new SSM layer H3 to fill the performance gap between SSMs and Transformers in language modeling. Mehta et al. [32] introduce a gated state space layer using gated units for improving expressivity. Recently, Gu & Dao [13] further propose Mamba with the core operation S6, an input-dependent selection mechanism of S4, which achieves linear scaling in sequence length and demonstrates superior performance over Transformers on various benchmarks. Mamba has been successfully applied in image classification [27, 54], image segmentation [41], and graph prediction [38]. Different from them, which use Mamba in the single-task setting, we consider a more challenging multi-task setting and propose novel self-task and cross-task Mamba modules to capture intra-task and inter-task dependence.
@@ 3.4 Mamba-based Decoder
Mamba-based Feature Extractor (MFE).
We introduce a Mamba-based feature extractor to learn the representation of 2D images. It is a critical module in the proposed Mamba-based decoder. As shown in Figure 2(a), motivated by [13], MFE consists of a linear layer used to expand the feature dimension by a controllable expansion factor $\alpha$, a convolution layer with an activation function for extracting local features, an SS2D operation for modeling long-range dependency, and a layer normalization to normalize the learned features. More formally, given the input feature ${\bf z}$, the output $\bar{{\bf z}}$ of MFE is calculated as
@@
$\tilde{{\bf z}}$ $=\texttt{MFE}({\bf z}_{\text{LN}}),$ (15)
${\bf g}$ $=\sigma(\texttt{Linear}({\bf z}_{\text{LN}})),$ (16)
@@ 4.1 Experimental Setups
Following [47, 43], experiments are conducted on two benchmark datasets with multi-task labels: NYUDv2 [35] and PASCAL-Context [6]. The NYUDv2 dataset comprises a variety of indoor scenes, containing 795 and 654 RGB images for training and testing, respectively. It consists of four tasks: $40$-class semantic segmentation (Semseg), monocular depth estimation (Depth), surface normal estimation (Normal), and object boundary detection (Boundary). The PASCAL-Context dataset, derived from the PASCAL dataset [9], includes both indoor and outdoor scenes and provides pixel-wise labels for tasks like semantic segmentation, human parsing (Parsing), and object boundary detection, with additional labels for surface normal estimation and saliency detection tasks generated by [31]. It contains 4,998 training images and 5,105 testing images.
@@
Table 1: Comparison with state-of-the-art methods on NYUDv2 (left) and PASCAL-Context (right) datasets. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance. The best and second best results are highlighted in bold and underline, respectively.
@@ 4.2 Comparison with State-of-the-art Methods
Table 1 shows the results on NYUDv2 and PASCAL-Context datasets. As can be seen, MTMamba shows superior performance in all four tasks on NYUDv2. For example, the performance of the semantic segmentation task has notably improved from the Transformer-based methods (i.e., InvPT and MQTransformer), increasing by +2.26 and +0.98, respectively, which demonstrates the effectiveness of MTMamba. The results on PASCAL-Context show the clear superiority of MTMamba. Especially, MTMamba significantly improves the previous best by +2.08, +5.01, and +4.90 in semantic segmentation, human parsing, and object boundary detection tasks, respectively, showing the effectiveness of MTMamba again. The qualitative comparison with InvPT on NYUDv2 and PASCAL-Context is shown in Figures 4 and 5, showing that MTMmaba provides more precise predictions and details.
Table 2: Effectiveness of the STM and CTM blocks on NYUDv2. Swin-Large encoder is used in this experiment. “Multi-task” denotes an MTL model where each task only uses two standard Swin Transformer blocks in each decoder stage. “Single-task” is the single-task counterpart of “Multi-task”. $\blacklozenge$, $\spadesuit$, $\blacksquare$, and $\bigstar$ are different variants of MTMamba. $\bigstar$ is the default configuration of MTMamba. ($\uparrow$($\downarrow$)↑ ( ↓ ) indicates that a higher (lower) result corresponds to better performance.
Method Each Decoder Semseg Depth Normal Boundary $\Delta_{m}$[%] #Param FLOPs
@@
Single-task 2*Swin 54.32 0.5166 19.21 77.30 0.00 888.77 1074.79
Multi-task 2*Swin 53.72 0.5239 19.97 76.50 -1.87 303.18 466.35
MTMamba 1*STM 54.61 0.5059 19.00 77.40 +0.95 252.51 354.13
2*STM 54.66 0.4984 18.81 78.20 +1.84 276.48 435.47
3*STM 54.75 0.5054 18.81 78.20 +1.55 300.45 516.82
2*STM+1*CTM 55.82 0.5066 18.63 78.70 +2.38 307.99 540.81
Table 3: Effectiveness of MFE module in MTMamba on NYUDv2. Swin-Large encoder is used. “W-MSA” is the window-based multi-head self-attention module in Swin Transformer [28]. “MFE” denotes all MFE modules in both STM and CTM blocks.
@@
Table 4: Effectiveness of linear gate in MTMamba on NYUDv2. Swin-Large encoder is used. “W-MSA” is the window-based multi-head self-attention module in Swin Transformer [28]. “Linear” denotes all linear gates in both STM and CTM blocks.
@@
4.3 Model Analysis
Effectiveness of STM and CTM Blocks.
The decoder of MTMamba contains two types of core blocks: STM and CTM blocks. We experiment on NYUDv2 to study the effectiveness of each type when the encoder is fixed as a Swin-Large Transformer. The results are shown in Table 2, where “Multi-task” represents an MTL model using two standard Swin Transformer blocks in each decoder stage for each task, and “Single-task” is the single-task counterpart of “Multi-task” (i.e., each task has a task-specific encoder-decoder). According to Table 2, the STM block achieves better performance and is more efficient than the Swin Transformer block ($\spadesuit$ vs. “Multi-task”), demonstrating that Mamba is more beneficial to multi-task dense prediction. Simply increasing the number of STM blocks from two to three fails to boost the performance. However, when the CTM is used, MTMamba has a significantly better performance in terms of $\Delta_{m}$ ($\bigstar$ vs. $\spadesuit$/$\blacksquare$). Moreover, the default configuration of MTMamba (i.e., $\bigstar$) significantly outperforms “Single-task” on all tasks, showing that MTMamba is more powerful.
@@ 4.3 Model Analysis
As shown in Figure 2, the MFE module is SSM-based and is the core of both STM and CTM blocks. We conduct an experiment by replacing all MFE modules in MTMamba with the attention module. As shown in Table 3, MFE is more effective and efficient than attention.
@@
The core of the CTM block is the cross-task interaction, i.e., Equation (24), where we fuse task-specific representation $\tilde{{\bf z}}^{t}$ and shared representation $\tilde{{\bf z}}^{\text{sh}}$ via a task-specific gate ${\bf g}^{t}$. In this experiment, we study its effectiveness by comparing it with the cases of ${\bf g}^{t}=0$ and ${\bf g}^{t}=1$. The experiments are conducted with a Swin-Large Transformer encoder on NYUDv2. The results are shown in Table 6. As can be seen, using a specific $\tilde{{\bf z}}^{t}$ (i.e., the case of ${\bf g}^{t}=0$) or shared $\tilde{{\bf z}}^{\text{sh}}$ (i.e., the case of ${\bf g}^{t}=1$) degrades the performance, demonstrating that the adaptive fusion is better.
@@ 4.4 Qualitative Evaluations
Figure 3 shows the comparison of the final decoder feature between MTMamba and Transformer-based method InvPT [47] in the semantic segmentation task. As can be seen, our method highly activates the regions with contextual and semantic information, which means it captures more discriminative features, resulting in better segmentation performance.
@@
We conduct qualitative studies by comparing the output predictions from our proposed MTMamba against the state-of-the-art Transformer-based method, InvPT [47]. Figures 4 and 5 show the qualitative results on the NYUDv2 and PASCAL-Context datasets, respectively. As can be seen, our method has better visual results than InvPT in all tasks. For example, as highlighted with yellow circles in Figure 4, MTMamba generates more accurate results with better alignments for the semantic segmentation task and clearer object boundaries for the object boundary detection task. Figure 5 demonstrates that MTMamba produces better predictions with more accurate details (like the fingers as highlighted) for both semantic segmentation and human parsing tasks and more distinct boundaries for the object boundary detection task. Hence, both qualitative study (Figures 4 and 5) and quantitative study (Table 1) show the superior performance of the proposed MTMamba method.
@@ 5 Conclusion
In this paper, we propose MTMamba, a novel multi-task architecture with a Mamba-based decoder for multi-task dense scene understanding. With two core blocks (STM and CTM blocks), MTMamba can effectively model long-range dependency and achieve cross-task interaction. Experiments on two benchmark datasets demonstrate that the proposed MTMamba achieves better performance than previous CNN-based and Transformer-based methods.
Acknowledgements
This work is supported by Guangzhou-HKUST(GZ) Joint Funding Scheme (No. 2024A03J0241).